The spring in a BB gun has a force constant of 1.8e4 N/m. When loaded, the spring is compressed a distance of 1.2cm. If the mass of the BB is 0.36g, then what is the speed of the BB as it is fired from the toy gun?
Answer #1

By Conservation of energy

(1/2)Kx2 = (1/2)mV2

(1.8*104)(0.012)2 = (0.36*10-3)V2

V2 =2.592/(0.36*10-3) =7200

V=84.85 m/s